what is the answer to the next problem if your solving for x and y using equal value method?
y=x-2
y=-x+2

Answer by stanbon(75887)

what is the answer to the next problem if your solving for x and y using equal value method?
y=x-2
y=-x+2
--------
Substitute and solve for "x":
x-2 = -x+2
2x = 4
x = 2
------
Solve for "y" using y = x-2
y = 2-2
y = 0
==========
Cheers,
Stan H.